OpenStudy (anonymous):

At Shaw's Market apples cost $10 per case and bananas cost $6 per case. If an order comes in for a total of 300 cases for $2,000, what was the specific number of cases of apples?

OpenStudy (anonymous):

10x+6(300-x)=2000. solve for x

OpenStudy (anonymous):

10x + 6(300-x) = 2000 => 4x + 1800 = 2000 => 4x = 200 => x = 50 (Ans)